Question:
FInd two consecutive even integers whose products is 168.
the equation i set up was n+(n+1)=168, is that right? i don't know what to do from there if it is.

Answer by scott8148(6628)

even (or odd) integers are 2 units apart, and product implies multiplication

n(n + 2) = 168 ___ n^2 + 2n - 168 = 0

factoring ___ (n + 14)(n - 12) = 0